Translate the following sentence into a mathematical equation. Be sure to identify the meaning of all symbols.
The profit derived from the sale of x video cameras is $370 per unit less the sum of $2700 costs plus $160 per unit.

Answer by Theo(13342)

the verbal statement is:

The profit derived from the sale of x video cameras is $370 per unit less the sum of $2700 costs plus $160 per unit.

what this statement is telling you is that the profit is equal to the revenue minus the cost.

the revenue is 370 per unit.
the cost is 2700 + 160 per unit.

the general formula is:

p = r - c

p is the profit.
r is the revenue
c is the cost.

r = 370 * x
c = 2700 + 160 * x

p = r - c becomes:

p = 370 * x - (2700 + 160 * x)

simplify to get:

p = 370 * x - 2700 - 160 * x

combine like terms to get:

p = 210 * x - 2700.